# Stochastic Control Methods for Optimization

### v1

@@ 3.1 Finite Particle Approximation
Proof.
Step 1. First, we estimate the first order derivative. For each $(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, $\xi\sim\mu$ under the above assumption, the stochastic Pontryagin maximum principle established in [11] implies that the optimal control $\theta^{*}$ exists, and the optimal state process $X^{t,\xi;\theta^{*}}$ satisfies the following FBSDE:
$\begin{cases}dX^{t,\xi;\theta^{*}}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s}\,ds+dW_{s},\quad X^{t,\xi;\theta^{*}}_{t}=\xi\sim\mu,\\ dY^{t,\xi;\theta^{*}}_{s}=Z^{t,\xi;\theta^{*}}_{s}\,dB_{s},\quad Y^{t,\xi;\theta^{*}}_{1}=\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}),\\ \theta^{*}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s},\quad\forall\,s\in[t,T].\end{cases}$ (3.1)
For any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})\text{ with }\xi_{1}\sim\mu_{1},\xi_{2}\sim\mu_{2}$, we denote by $(X^{t,\xi_{i};\theta^{*}},Y^{t,\xi_{i};\theta^{*}},Z^{t,\xi_{i};\theta^{*}})$ the solution to the FBSDE (3.1) with initial condition $\xi_{i}$, $i=1,2$. Put $\Delta X_{s}=X^{t,\xi_{1};\theta^{*}}_{s}-X^{t,\xi_{2};\theta^{*}}_{s}$, $\Delta Y_{s}=Y^{t,\xi_{1};\theta^{*}}_{s}-Y^{t,\xi_{2};\theta^{*}}_{s}$, and $\Delta Z_{s}=Z^{t,\xi_{1};\theta^{*}}_{s}-Z^{t,\xi_{2};\theta^{*}}_{s}$. Applying Itô’s formula to $\langle\Delta X_{s},\Delta Y_{s}\rangle$ and taking expectations yield that
$\mathbb{E}\left[\langle\Delta X_{1},\Delta Y_{1}\rangle\right]$ $=-\frac{1}{\varepsilon}\mathbb{E}\left[\int_{r}^{1}|\Delta Y_{s}|^{2}ds\right]+\mathbb{E}\left[\langle\Delta Y_{r},\,\Delta X_{r}\rangle\right],\forall\,r\in[t,1].$ (3.2)
On the other hand, observing that the convexity implies the monotonicity of $\partial_{\mu}G$, we have
$\mathbb{E}\left[\langle\Delta X_{1},\Delta Y_{1}\rangle\right]=\mathbb{E}\left[\langle X^{t,\xi_{1};\theta^{*}}_{1}-X^{t,\xi_{2};\theta^{*}}_{1},\partial_{\mu}G(\mathcal{L}(X^{t,\xi_{1};\theta^{*}}_{1}))(X^{t,\xi_{1};\theta^{*}}_{1})-\partial_{\mu}G(\mathcal{L}(X^{t,\xi_{2};\theta^{*}}_{1}))(X^{t,\xi_{2};\theta^{*}}_{1})\rangle\right]\geq 0.$
Combining the above two estimates yields that
$0\leq\mathbb{E}\left[\langle\Delta Y_{r},\,\Delta X_{r}\rangle\right],\forall\,r\in[t,1].$ (3.3)
Again, Applying Itô’s formula to $|\Delta X_{r}|^{2}$ and taking expectations yield that $\mathbb{E}[|\Delta X_{r}|^{2}]\leq\mathbb{E}[|\xi_{1}-\xi_{2}|^{2}],\forall\,r\in[t,1].$ Thus, in view of the BSDE and Lipschitz continuity of $\partial_{\mu}G$ due to the boundedness of the second order derivatives $\partial^{2}_{\mu\mu}G$ and $\partial_{x}\partial_{\mu}G$, we have
$\sup_{s\in[t,T]}\mathbb{E}\left[|\Delta Y_{s}|\right]\leq\sup_{s\in[t,T]}\sqrt{\mathbb{E}\left[|\Delta Y_{s}|^{2}\right]}$ $\leq 2K_{G}\sqrt{\mathbb{E}\left[|\Delta X_{1}|^{2}\right]}\leq 2K_{G}\sqrt{\mathbb{E}[|\xi_{1}-\xi_{2}|^{2}]},$ (3.4)
which implies the continuity of the optimal control $\theta^{*}$ in $\mu$.
Let $\phi\in L^{2}(\Omega;\mathbb{R}^{d})$ be a perturbation of the initial position $\xi$. Let $\xi^{\delta}=\xi+\delta\phi$ for $\delta\in(0,1)$. Let $\mu^{\delta}=\mathcal{L}(\xi^{\delta})$ and $\theta^{\delta}$ be the optimal control associated to $(t,\xi^{\delta})$. Then it is obvious that
$\frac{J_{\varepsilon}^{\theta^{\delta}}(t,\mu^{\delta})-J_{\varepsilon}^{\theta^{\delta}}(t,\mu)}{\delta}\leq\frac{\mathcal{V}_{\varepsilon}(t,\mu^{\delta})-\mathcal{V}(t,\mu)}{\delta}\leq\frac{J_{\varepsilon}^{\theta^{*}}(t,\mu^{\delta})-J_{\varepsilon}^{\theta^{*}}(t,\mu)}{\delta}$
with
$\frac{J_{\varepsilon}^{\theta^{*}}(t,\mu^{\delta})-J_{\varepsilon}^{\theta^{*}}(t,\mu)}{\delta}=\frac{\mathbb{E}[G(\mathcal{L}(X^{t,\xi^{\delta};\theta^{*}}_{1}))]-\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))]}{\delta}=\frac{\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}+\delta\phi))]-\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))]}{\delta},$
$\frac{J_{\varepsilon}^{\theta^{\delta}}(t,\mu^{\delta})-J_{\varepsilon}^{\theta^{\delta}}(t,\mu)}{\delta}=\frac{\mathbb{E}[G(\mathcal{L}(X^{t,\xi^{\delta};\theta^{\delta}}_{1}))]-\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{\delta}}_{1}))]}{\delta}=\frac{\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{\delta}}_{1}+\delta\phi))]-\mathbb{E}[G(\mathcal{L}(X^{t,\xi;\theta^{\delta}}_{1}))]}{\delta}.$
Letting $\delta\to 0$ and using the differentiability of $G$ and the continuity of the optimal control with respect to the initial condition, we obtain by the Squeeze Theorem that
$\lim_{\delta\to 0}\frac{\mathcal{V}_{\varepsilon}(t,\mu^{\delta})-\mathcal{V}(t,\mu)}{\delta}=\mathbb{E}\left[\left\langle\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}),\phi\right\rangle\right]=\mathbb{E}[Y_{t}^{t,\xi;\theta^{*}}\cdot\phi].$
Thus, $\mathcal{V}_{\varepsilon}$ is Lions differentiable with
$\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu,x)=Y_{t}^{t,x;\theta^{*}}.$ (3.5)
Step 2. Next, we estimate the second order derivatives. It becomes clear that to study the higher-order regularity of $\mathcal{V}_{\varepsilon}$, we need to differentiate the FBSDE (3.1) with respect to the initial condition $x$ and the measure $\mu$. Let $\xi\sim\mu$. As the calculations are standard, we sketch the main steps and refer to [10, 12, 16] for details.
We first differentiate the FBSDE (3.1) with respect to the measure $\mu$. Let $\phi\in L^{2}(\mathbb{R}^{d},\mu;\mathbb{R}^{d})$ be a vector field representing the direction of the perturbation. This induces a perturbation of the initial random variable $\xi$ given by $\delta X^{\phi}_{t}=\phi(\xi)\in L^{2}(\Omega;\mathbb{R}^{d})$. The tangent process $(\delta X_{s},\delta Y_{s},\delta Z_{s})$ is defined as the $L^{2}(\Omega)$-limit of the difference quotients:
where $\theta^{r}$ is the optimal control associated to the initial condition $\xi+r\phi(\xi)$. Differentiating the FBSDE coefficients leads to the following linearized system (variational equation):
$\begin{cases}d(\delta X_{s})=-\frac{1}{\varepsilon}\delta Y_{s}\,ds,\quad\delta X_{t}=\phi(\xi),\\ d(\delta Y_{s})=\delta Z_{s}\,dB_{s},\\ \delta Y_{1}=\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1},\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}].\end{cases}$ (3.6)
In a similar manner to (3.2)-(3.3), applying Itô’s formulas, we obtain $|\delta X_{1}|\leq|\phi(\xi)|$ a.s. and further,
$\mathbb{E}[|\delta Y_{t}|]$ $=\mathbb{E}\left[\left|\mathbb{E}\left[\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1},\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}]\big|\mathcal{F}_{t}\right]\right|\right]$
$\leq K_{G}\mathbb{E}[|\delta X_{1}|]+K_{G}\mathbb{E}[|\delta\tilde{X}_{1}|]$
$\leq 2K_{G}\sqrt{\mathbb{E}[|\phi(\xi)|^{2}]},$
which by the arbitrariness of $\phi$ implies that $|\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x,y)|\leq 2K_{G}$ for all $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$.
Analogously, following the standard linearization arguments for McKean–Vlasov FBSDEs (see, e.g., [12, 10, 16]), and differentiating the FBSDE (3.1) with respect to the spatial initial condition $x$, we may obtain the FBSDE for the spatial tangent processes $\big(\partial_{x}X^{t,x;\theta^{*}}_{s},\ \partial_{x}Y^{t,x;\theta^{*}}_{s},\ \partial_{x}Z^{t,x;\theta^{*}}_{s}\big)_{s\in[t,1]}$ to fetch the estimate. Indeed, the Lipschitz continuity in (3.4) implies that
$|\partial_{x}Y^{t,x;\theta^{*}}_{t}|=|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|\leq 2K_{G}.$
Further, through standard stability analysis of linear BSDEs (see e.g. [12]), we may obtain the continuity of the maps $(t,x,\mu)\mapsto\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)$ and $(t,\mu)\mapsto\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)$.
Step 3. We are ready to show that $\mathcal{V}_{\varepsilon}$ is the unique classical solution to the Master Equation (1.20). For each time $t\in[0,1]$, we have proved that the optimal control $\theta^{*}_{t}=-\frac{1}{\varepsilon}Y$ as a function of $(\mu,x)$ is uniformly Lipschitz continuous and applying standard BSDE techniques, we have the joint continuous in $(t,x,\mu)$ (see [11, Section 5.6]). Thus, we may restrict the admissible controls to feedback controls of the form $\theta_{s}=\alpha(s,\mathcal{L}(X_{s}^{t,\xi;\theta}),X_{s}^{t,\xi;\theta})$ with $\alpha:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ Lipschitz continuous (denote by $\hat{L}$ the Lipschitz constant) in $(\mu,x)$ uniformly in time and $\sup_{t\in[0,1]}|\alpha(t,\delta_{0},0)|\leq M$ for some constant $M>0$. And accordingly, using Itô formula, we get the uniform estimates of the state processes, i.e., $\mathbb{E}[|X_{s}^{t,\xi;\alpha}|^{2}]\leq C(\hat{L},M)\mathbb{E}[|\xi|^{2}+1]$ with $C(\hat{L},M)$ independent of $(s,t,\xi)$, and thus, conversely, we obtain the boundedness of the such controls, i.e.,
$\mathbb{E}[|\alpha(s,\mathcal{L}(X_{s}^{t,\xi;\alpha}),X_{s}^{t,\xi;\alpha})|^{2}]\leq\tilde{C}(\hat{L},M)(1+\mathbb{E}[|\xi|^{2}]),\quad\forall\,s\in[t,T].$ (3.7)
For each $t\in[0,T]$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let $\xi\sim\mu$ and $\Theta(t,\mu):=\{\theta\in\Theta|\,\mathbb{E}[|\theta_{s}|^{2}]\leq\tilde{C}(\hat{L},M)(1+\mathbb{E}[|\xi|^{2}]),\,\,\forall\,s\in[t,T]\}$. Then, for any $t,s\in[0,1]$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we have
$|\mathcal{V}_{\varepsilon}(t,\mu)-\mathcal{V}_{\varepsilon}(s,\mu)|\leq\sup_{\theta\in\Theta(t\wedge s,\mu)}\mathbb{E}\left[\int_{s\wedge t}^{t\vee s}|\theta_{r}|^{2}dr\right]\leq\tilde{C}(\hat{L},M)(1+\mathbb{E}[|\xi|^{2}])|t-s|,\quad\text{for }\xi\sim\mu.$
By Rademacher’s theorem, $\mathcal{V}_{\varepsilon}$ is differentiable in time almost everywhere ($dt$-a.e.).
Applying generalized Itô-Kunita-Wentzell formula for functions on Wasserstein space (see e.g. [24]), we have for $s\in[t,T]$:
$\mathcal{V}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))$ $=\mathcal{V}_{\varepsilon}(t,\mu)+\int_{t}^{s}\partial_{t}\mathcal{V}_{\varepsilon}(r,\mathcal{L}(X^{t,\xi;\theta^{*}}_{r}))dr-\frac{1}{\varepsilon}\int_{t}^{s}\mathbb{E}\left[\left|\partial_{\mu}\mathcal{V}_{\varepsilon}(r,\mathcal{L}(X^{t,\xi;\theta^{*}}_{r}),X^{t,\xi;\theta^{*}}_{r})\right|^{2}\right]dr$
$\quad+\frac{1}{2}\int_{t}^{s}\mathbb{E}\left[\text{div}_{x}(\partial_{\mu}\mathcal{V}_{\varepsilon})(r,\mathcal{L}(X^{t,\xi;\theta^{*}}_{r}),X^{t,\xi;\theta^{*}}_{r})\right]dr.$ (3.8)
By the dynamic programming principle (see [12]), it holds that
$\mathcal{V}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))+\int_{t}^{s}\frac{1}{2\varepsilon}\mathbb{E}\left[\left|\partial_{\mu}\mathcal{V}_{\varepsilon}(r,\mathcal{L}(X^{t,\xi;\theta^{*}}_{r}),X^{t,\xi;\theta^{*}}_{r})\right|^{2}\right]dr=\mathcal{V}_{\varepsilon}(t,\mu).$
Substituting the above equality into (3.8) and after cancellations taking the derivative with respect to $s$ at $s=t$ yields that $\mathcal{V}_{\varepsilon}$ admits a continuous time-derivative and is a classical solution to the Master Equation (1.20). Further by Lemma 3.1, the solution is unique. ∎
@@
Let Assumption 3.1 hold and let $v^{N}_{\varepsilon}(t,\mathbf{x})$ be the value function of the $N$-particle control problem satisfying the HJB equation (1.23). Then, for any configuration $\mathbf{x}\in\mathbb{R}^{Nd}$, the per-particle value function converges with rate $O(1/N)$:
@@
The proof follows an approach as used in [31] relies on projecting the Master Equation solution onto the finite-dimensional manifold and using the comparison principle for the HJB equation, is provided in Appendix A.
@@
$|\mathcal{V}_{\varepsilon}(t,\mu_{1})-\mathcal{V}_{\varepsilon}(t,\mu_{2})|$ $=\Big|\inf_{\theta\in\Theta}J^{\theta}_{\varepsilon}(t,\mu_{1})-\inf_{\theta\in\Theta}J^{\theta}_{\varepsilon}(t,\mu_{2})\Big|$
$\leq\sup_{\theta\in\Theta}\left|G(\mathcal{L}(X_{1}^{t,\mu_{1};\theta}))-G(\mathcal{L}(X_{1}^{t,\mu_{2};\theta}))\right|$
@@
$\mathbb{E}[\mathcal{W}_{2}(\mu^{N}_{\mathbf{X}},\mu)]\leq\begin{cases}C_{\alpha}N^{-1/2-\alpha},\quad\forall\,\alpha>0,&\text{if }d\leq 4,\\ C_{2}N^{-1/d},&\text{if }d>4,\end{cases}$
@@ 3.2 Regularization error
Let $G:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be Lipschitz continuous with respect to the Wasserstein-2 distance with constant $L_{G}$. Assume that $G$ admits a global minimizer $\mu^{*}$ with a finite second moment (i.e., $\int|x|^{2}d\mu^{*}(x)<\infty$). Let $\mathcal{V}_{0}=\inf_{\mu}G(\mu)$. Then the value function $\mathcal{V}_{\varepsilon}$ satisfies:
$0\leq\mathcal{V}_{\varepsilon}-\mathcal{V}_{0}\leq C\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right),$ (3.11)
for some constant $C>0$ depending on the dimension $d$, the Lipschitz constant $L_{G}$, and the Wasserstein-2 distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between the Dirac measure at $x_{0}$ and $\mu^{*}$, as $\varepsilon\to 0$.
@@
Step 1. First, we reformulate the stochastic control problem as a variational problem over probability measures. Let $\mathbb{P}$ be the Wiener measure (the law of the uncontrolled process where $\theta_{t}\equiv 0$). For any admissible control $\theta\in\Theta$, let $\mathbb{Q}^{\theta}$ be the law of the controlled process $X$.
By Girsanov’s theorem, the Radon-Nikodym derivative of $\mathbb{Q}^{\theta}$ with respect to $\mathbb{P}$ is:
@@
The relative entropy (Kullback-Leibler divergence) is given by:
@@
This identity allows us to rewrite the cost functional purely in terms of the induced measure $\mathbb{Q}^{\theta}$, and the control problem becomes
$\mathcal{V}_{\varepsilon}=\inf_{\theta\in\Theta}\left\{G(\mathcal{L}(X_{1}^{\mathbb{Q}^{\theta}}))+\varepsilon\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\right\}.$ (3.12)
The lower bound follows immediately from the non-negativity of the relative entropy term, i.e., as for any admissible control $\theta$, $\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\geq 0$, we have
$\mathcal{V}_{\varepsilon}=\inf_{\theta}\left\{G(\mathcal{L}(X_{1}^{\mathbb{Q}^{\theta}}))+\varepsilon\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\right\}\geq\inf_{\theta}G(\mathcal{L}(X_{1}^{\mathbb{Q}^{\theta}}))\geq\inf_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}G(\mu)=\mathcal{V}_{0}.$ (3.13)
Step 2. Next, we prove the upper bound by constructing a specific suboptimal control strategy $\theta^{\delta}$. Let $\mu^{*}$ be a global minimizer of $G$, i.e., $G(\mu^{*})=\mathcal{V}_{0}$. By assumption, $\mu^{*}$ has a finite second moment. Note that $\mu^{*}$ may be singular (e.g., a Dirac measure). To handle this, we smooth $\mu^{*}$ by convolving it with a Gaussian kernel. We define a smoothed candidate measure $\mu^{\delta}$ by convolving $\mu^{*}$ with a Gaussian kernel $\phi_{\delta}$ of variance $\delta\in(0,1)$, i.e.,
$\mu^{\delta}(dx)=(\mu^{*}\ast\phi_{\delta})(x)dx=\int_{\mathbb{R}^{d}}\frac{1}{(2\pi\delta)^{d/2}}\exp\left(-\frac{\|x-y\|^{2}}{2\delta}\right)\,\mu^{*}(dy)\,dx.$ (3.14)
The theory of Schrödinger bridges (see, e.g., [28, 45]) guarantees the existence of such a control $\theta^{\delta}$ that the law of the controlled process at time $t=1$ matches $\mu^{\delta}$, i.e., $\mathcal{L}(X_{1}^{\theta^{\delta}})=\mu^{\delta}$. Moreover, the minimum entropy cost to achieve a marginal distribution $\mu^{\delta}$ at time $t=1$ is simply the relative entropy of the marginals:
$\inf_{\theta:\mathcal{L}(X_{1}^{\theta})=\mu^{\delta}}\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})=\mathcal{H}(\mu^{\delta}|\gamma_{x_{0}}),$
where $\gamma_{x_{0}}=\mathcal{L}(W_{1}+x_{0})\sim\mathcal{N}(x_{0},I_{d})$. Substituting this control into the cost functional gives an upper bound:
$\mathcal{V}_{\varepsilon}\leq G(\mu^{\delta})+\varepsilon\mathcal{H}(\mu^{\delta}|\gamma_{x_{0}}).$ (3.15)
As $G$ is $L_{G}$-Lipschitz with respect to $\mathcal{W}_{2}$, we have $|G(\mu^{\delta})-G(\mu^{*})|\leq L_{G}\cdot\mathcal{W}_{2}(\mu^{\delta},\mu^{*}).$ Recalling that the measure $\mu^{\delta}$ is obtained by adding independent noise $\sqrt{\delta}Z$ to $\mu^{*}$, with $Z\sim\mathcal{N}(0,I_{d})$. The Wasserstein-2 distance is bounded by the second moment of the noise. i.e.,
$\mathcal{W}_{2}(\mu^{\delta},\mu^{*})\leq\sqrt{\mathbb{E}[|\sqrt{\delta}Z|^{2}]}=\sqrt{d\delta}.$
@@
$G(\mu^{\delta})\leq\mathcal{V}_{0}+L_{G}\sqrt{d\delta}.$

### v2

@@ 3.1 Finite Particle Approximation
The identification of the value function $\mathcal{V}_{\varepsilon}$ as a classical solution to the master equation (1.20) follows from the stochastic Pontryagin maximum principle in [12] and the classical solution theory of master equations in [16]; in particular, this justification does not require the assumption (c) (Coercivity) in Proposition 3.2. In contrast, the coercivity condition plays an essential role in our argument for deriving the $\varepsilon$-uniform bounds in (3.1). Our analysis is devoted to establishing these $\varepsilon$-uniform estimates, which, to the best of our knowledge, are not addressed explicitly in the existing literature. For the sake of readability, the technical proof is deferred to Appendix A.
@@
Let Assumption 3.1 hold and let $v^{N}_{\varepsilon}(t,\mathbf{x})$ be the value function of the $N$-particle control problem satisfying the HJB equation (1.23). Then, for any configuration $\mathbf{x}\in\mathbb{R}^{Nd}$, the per-particle value function converges with rate $O(1/N)$, i.e.,
@@
A proof, inspired by the projection argument developed in [30] and based on combining the projection of the master-equation solution onto the empirical manifold with a comparison principle for the finite-dimensional HJB equation, is given in Appendix B.
@@
$|\mathcal{V}_{\varepsilon}(t,\mu_{1})-\mathcal{V}_{\varepsilon}(t,\mu_{2})|=\Big|\inf_{\theta\in\Theta}J^{\theta}_{\varepsilon}(t,\mu_{1})-\inf_{\theta\in\Theta}J^{\theta}_{\varepsilon}(t,\mu_{2})\Big|$ $\leq\sup_{\theta\in\Theta}\left|G(\mathcal{L}(X_{1}^{t,\mu_{1};\theta}))-G(\mathcal{L}(X_{1}^{t,\mu_{2};\theta}))\right|$
@@
$\mathbb{E}[\mathcal{W}_{2}(\mu^{N}_{\mathbf{X}},\mu)]\leq\begin{cases}C_{\alpha}N^{-1/2+\alpha},\quad\text{for}\,\alpha\in(0,\frac{1}{2}),&\text{if }d\leq 4,\\ C_{2}N^{-1/d},&\text{if }d>4,\end{cases}$
@@ 3.2 Regularization error
Let $G:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be Hölder continuous with respect to the Wasserstein-2 distance, i.e., there exist $L_{G}>0$ and $\alpha\in(0,1]$ such that $|G(\mu)-G(\nu)|\leq L_{G}|\mathcal{W}_{2}(\mu,\nu)|^{\alpha}$ for all $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Assume that $G$ admits a global minimizer $\mu^{*}$ with a finite second moment (i.e., $\int|x|^{2}d\mu^{*}(x)<\infty$). Let $\mathcal{V}_{0}=\inf_{\mu}G(\mu)$. Then the value function $\mathcal{V}_{\varepsilon}$ satisfies
$0\leq\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})-\mathcal{V}_{0}\leq C\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right),\quad\text{for }\varepsilon\in(0,e^{-1}),$ (3.4)
for some constant $C>0$ depending on $d$, $L_{G}$, $\alpha$, and the Wasserstein-2 distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between the Dirac measure at $x_{0}$ and $\mu^{*}$.
@@
Step 1. First, we reformulate the stochastic control problem as a variational problem over probability measures. Let $\mathbb{P}$ be the Wiener measure for paths starting at $x_{0}$ at time 0 (the law of the uncontrolled proces $X^{0,x_{0};0}$ where $\theta\equiv 0$). For any admissible control $\theta\in\Theta$, let $\mathbb{Q}^{\theta}$ be the law of the controlled process $X^{0,x_{0};\theta}$.
By Girsanov’s theorem, the Radon-Nikodym derivative of $\mathbb{Q}^{\theta}$ with respect to $\mathbb{P}$ is
@@
The relative entropy (Kullback-Leibler divergence) is given by
@@
This identity allows us to rewrite the cost functional in terms of the induced measure $\mathbb{Q}^{\theta}$, and the control problem becomes
$\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})=\inf_{\theta\in\Theta}\left\{G(\mathcal{L}(X_{1}^{0,x_{0};\theta}))+\varepsilon\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\right\}.$ (3.5)
The lower bound is obvious from the control problem and alternatively, it follows immediately from the non-negativity of the relative entropy term, i.e., as for any admissible control $\theta$, $\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\geq 0$, we have
$\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})=\inf_{\theta}\left\{G(\mathcal{L}(X_{1}^{0,x_{0};\theta}))+\varepsilon\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})\right\}\geq\inf_{\theta}G(\mathcal{L}(X_{1}^{0,x_{0};\theta}))\geq\inf_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}G(\mu)=\mathcal{V}_{0}.$ (3.6)
Step 2. Next, we prove the upper bound by constructing a specific suboptimal control strategy $\theta^{\lambda}$. Let $\mu^{*}$ be a global minimizer of $G$, i.e., $G(\mu^{*})=\mathcal{V}_{0}$. By assumption, $\mu^{*}$ has a finite second moment. Note that $\mu^{*}$ may be singular (e.g., a Dirac measure). To handle this, we smooth $\mu^{*}$ by convolving it with a Gaussian kernel. We define a smoothed candidate measure $\mu^{\lambda}$ by convolving $\mu^{*}$ with a Gaussian kernel $\phi_{\lambda}$ of variance $\lambda\in(0,1)$, i.e.,
$\mu^{\lambda}(dx)=(\mu^{*}\ast\phi_{\lambda})(x)dx=\int_{\mathbb{R}^{d}}\frac{1}{(2\pi\lambda)^{d/2}}\exp\left(-\frac{|x-y|^{2}}{2\lambda}\right)\,\mu^{*}(dy)\,dx.$ (3.7)
The theory of Schrödinger bridges (see, e.g., [27, 44]) guarantees the existence of such a control $\theta^{\lambda}$ that the law of the controlled process at time $t=1$ matches $\mu^{\lambda}$, i.e., $\mathcal{L}(X_{1}^{0,x_{0};\theta^{\lambda}})=\mu^{\lambda}$. Moreover, the minimum entropy cost to achieve a marginal distribution $\mu^{\lambda}$ at time $t=1$ is simply the relative entropy of the marginals, i.e.,
$\inf_{\theta:\mathcal{L}(X_{1}^{0,x_{0};\theta})=\mu^{\lambda}}\mathcal{H}(\mathbb{Q}^{\theta}|\mathbb{P})=\mathcal{H}(\mu^{\lambda}|\gamma_{x_{0}}),$
where $\gamma_{x_{0}}=\mathcal{L}(W_{1}+x_{0})\sim\mathcal{N}(x_{0},I_{d})$. Substituting this control into the cost functional gives an upper bound
$\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})\leq G(\mu^{\lambda})+\varepsilon\mathcal{H}(\mu^{\lambda}|\gamma_{x_{0}}).$ (3.8)
As $G$ is Hölder continuous with respect to $\mathcal{W}_{2}$, we have $|G(\mu^{\lambda})-G(\mu^{*})|\leq L_{G}\cdot|\mathcal{W}_{2}(\mu^{\lambda},\mu^{*})|^{\alpha}.$ Recalling that the measure $\mu^{\lambda}$ is obtained by adding independent noise $\sqrt{\lambda}Z$ to $\mu^{*}$, with $Z\sim\mathcal{N}(0,I_{d})$. The Wasserstein-2 distance is bounded by the second moment of the noise. i.e.,
$\mathcal{W}_{2}(\mu^{\lambda},\mu^{*})\leq\sqrt{\mathbb{E}[|\sqrt{\lambda}Z|^{2}]}=\sqrt{d\lambda}.$
@@
$G(\mu^{\lambda})\leq\mathcal{V}_{0}+L_{G}(d\lambda)^{\alpha/2}.$